# Two-dimensional fluids via matrix hydrodynamics

### v2

@@ Abstract.
Two-dimensional (2-D) incompressible, inviscid fluids produce fascinating patterns of swirling motion. How and why the patterns emerge are long-standing questions, first addressed in the 19th century by Helmholtz, Kirchhoff, and Kelvin. Countless researchers have since contributed to innovative techniques and results, but the overarching problem of swirling 2-D motion and its long-time behavior remains largely open. Here we advocate an alternative view-point that sheds light on this problem via a link to isospectral matrix flows. The link is established through V. Zeitlin’s beautiful model for the numerical discretization of Euler’s equations in 2-D. When considered on the sphere, Zeitlin’s model enables a deep connection between 2-D hydrodynamics and unitary representation theory of Lie algebras as pursued in quantum theory. Consequently, it provides a dictionary that maps hydrodynamical concepts to matrix Lie theory, which in turn gives connections to matrix factorizations, random matrices, and integrability theory, for example. The transferal of outcomes, from finite-dimensional matrices to infinite-dimensional fluids, is then supported by hard-fought results in quantization theory – the field which describes the limit between quantum and classical physics. We demonstrate how the dictionary is constructed and how it unveils techniques for 2-D hydrodynamics. We also give accompanying convergence results for Zeitlin’s model on the sphere.
@@
Acknowledgements. The authors would like to thank Michael Roop for insightful comments, particularly related to the proof of Theorem 4, and Michele Dolce and Theo Drivas for fruitful discussions related to vortex mixing.
@@ 1. Introduction
If you stir up an incompressible, low-viscosity fluid confined to a thin (almost 2-D) bounded domain, what you see is spectacular, completely different from the situation in full 3-D. The fluid self-organizes into large, coherently swirling vortex condensates. Some regions swirl counterclockwise (positive vorticity) while others swirl clockwise (negative vorticity), with occasional merger between equal-signed regions, and repulsion between opposite-signed regions. What is the mechanism of vortex condensation and under which conditions does it occur? More mathematically formulated, what is the generic long-time behavior of the 2-D Euler equations?
@@
A milestone was reached in 1949, when Onsager [33] suggested to use statistical mechanics applied to a large but finite number $N$ of point vortices on a flat torus (doubly periodic square). Onsager ingeniously realized that if the energy of a configuration is large enough, the thermodynamical temperature is negative. Point vortices of equal sign will then tend to cluster “so as to use up excess energy at the least possible cost in terms of degrees of freedom” [33, p. 281]. Thus, statistical mechanics predicts vortex condensation. Under the ergodicity assumption, it also predicts the long-time states corresponding to “thermal equilibrium”. Onsager’s ideas have fostered several mathematically rigorous results, for example weak convergence, in both positive and negative temperature regimes, of the canonical Gibbs measure as $N\to\infty$ [2, 22]. However, there is a problem with the point vortex approach, as Onsager himself pointed out: “When we compare our idealized model with reality, we have to admit one profound difference: the distribution of vorticity which occur in the actual flow of normal liquids are continuous” [33, p. 281]. In particular, solutions to the 2-D Euler equations with continuous vorticity (more precisely $L^{\infty}$) possess a rich geometric structure: they are described by an infinite-dimensional Lie–Poisson system, which gives rise to infinitely many conservation laws given by Casimir functions (cf. Sec. 2 below). These Casimirs are not preserved (not even defined) in the point vortex model. Yet, there is evidence that the Casimirs significantly affect the long-time behavior (cf. Abramov and Majda [1]). From this perspective, Onsager’s model is not satisfactory. Furthermore, the ergodicity assumption is not valid (which does not, however, wholly invalidate the theory; see the discussion in Marchioro and Pulvirenti [28, Sec. 7.5]).
@@
Another established technique for 2-D hydrodynamics is non-linear PDE analysis. It has provided a flora of rigorous and deep results, but at the cost of somewhat restricted settings, for example perturbations of steady solutions. We refer to the monograph of Marchioro and Pulvirenti [28] and the lecture notes of Šverák [36] for an overview. The grand vision, of course, is to bridge PDE analysis results with predictions from statistical hydrodynamics. It is a notoriously difficult problem. A flavor of the difficulty is captured in the following “incompatibility” between the two approaches: The statistically predicted vortex condensation implies merging of equally signed vorticity regions. Yet, in the 2-D Euler equations, vorticity is transported by the fluid velocity field, so merging can only occur by trapping increasingly narrow, but deep, vorticity variations in an overwhelmingly complicated stretch-and-fold process. These variations cannot “disappear” in the $C^{1}$-topology required for the strong formulation of the governing PDE. Indeed, the $C^{1}$-norm along regular solutions can, and in numerical simulations typically does, grow extremely fast (e.g., double exponential [24]).
@@
A third tool for 2-D hydrodynamics is to carry out numerical experiments (e.g., Fig. 1, where the mixing phase (b) captures a typical stretch-and-fold process). Indeed, since the work of Lorentz [27] in the 1960s, computer simulations have guided many analytical results in fluid dynamics; recent examples are found in the references [6, 35, 24, 10, 4, 38]. But how can we know if the numerical results are consistent with the geometric structures principal to the long-time behavior? The key is to ensure that the finite-dimensional dynamical model resulting from the discretization is itself a Lie–Poisson system. Such a model was provided by Zeitlin [40, 41], based on quantization results of Hoppe [16, 17]. It is the only known self-consistent spatial discretization of 2-D Euler, yet surprisingly underexplored.
@@
Our paper draws from the following observation: in addition to numerical advantages, matrix hydrodynamics – the isospectral matrix flows enabled via Zeitlin’s model – brings theoretical insights to classical 2-D hydrodynamics, as it provides a link to matrix theory with all its findings. We therefore advocate matrix hydrodynamics as a tool for 2-D turbulence, similar in spirit to the point vortex model of Onsager, but with the benefit that it captures the Lie–Poisson geometry of continuous vorticity fields, including conservation of Casimirs. In addition to the examples in this paper, a motivation and feasibility check for the approach comes from complex geometry. Indeed, in the late 1980s, quantization was promoted to address problems related to Kähler–Einstein metrics and the Calabi conjecture. Tian [37] then gave convergence results, since used regularly in theoretical contexts. Only much later did Donaldson [8] use quantization in a numerical study of Kähler–Einstein metrics. This is the antithesis to the developments in 2-D hydrodynamics, where Zeitlin’s model has been a numerical tool, until recently.
@@
The primary aim of the paper is to give basic convergence results endorsing matrix hydrodynamics as an approach (Theorem 4, Theorem 6, Corollary 14, and Theorem 16), and to showcase the notion in selected examples and relate it to existing theories (Sec. 4-6). The secondary aim, addressed in Sec. 3, is to give a transparent derivation of Zeitlin’s model on the sphere, based on unitary representation theory and focused on geometry rather than algebraic formulae. Before that, we review in Sec. 2 the full geometric structure of the 2-D Euler equations.
@@ 2. Background: geometry of the 2-D Euler equations
Consider a domain filled with an infinite number of fluid particles that push each other in a frictionless, incompressible way. For such systems, Euler [11] showed how Newton’s equations give rise to a set of partial differential equations (PDEs) for the velocity field $\mathbf{v}(x,t)$ of the particles’ motion,
$\frac{\partial\mathbf{v}}{\partial t}+\nabla_{\mathbf{v}}\mathbf{v}=-\nabla p,\quad\operatorname{div}\mathbf{v}=0,$ (1)
where $p$ is the pressure function and $\nabla_{\mathbf{v}}$ is the co-variant derivative. Euler’s equations (1) make sense on any Riemannian manifold, but in 2-D the geometric structure is richer than in higher dimensions, and the dynamics is significantly different. Throughout this paper we shall work on the sphere $\mathbb{S}^{2}$, partly because of its relevance in geophysical contexts, but mostly because it later allows a description in terms of unitary representation theory that renders the geometric structures in Zeitlin’s model more transparent.
Let $J\colon T\mathbb{S}^{2}\to T\mathbb{S}^{2}$ denote the bundle mapping for rotation by $\pi/2$ in positive direction. The curl operator maps the vector field $\mathbf{v}$ on $\mathbb{S}^{2}$ to the vorticity function
$\omega=\operatorname{curl}\mathbf{v}\equiv\operatorname{div}(J\mathbf{v}).$
@@
$\frac{\partial\omega}{\partial t}+\operatorname{div}(\omega\mathbf{v})=0.$
Further, since the first co-homology of the sphere is trivial, it follows from Hodge theory that the divergence free vector field $\mathbf{v}$ can be written $\mathbf{v}=\nabla^{\bot}\psi\equiv-J\nabla\psi$, where $\psi$ is the stream function, unique up to a constant. Euler’s equations (1) can now be written entirely in terms of the vorticity and stream functions
$\frac{\partial\omega}{\partial t}+\operatorname{div}(\omega\nabla^{\bot}\psi)=0,\quad\Delta\psi=\omega.$ (2)
We now demonstrate how to arrive at these equations from the point-of-view of symplectic geometry and Hamiltonian dynamics.
The sphere is a symplectic manifold for the symplectic form given by the spherical area form $\mu$ (in spherical coordinates $(\theta,\phi)\in[0,\pi]\times[0,2\pi)$ the area form is given by $\mu=\sin\theta\,\mathrm{d}\theta\wedge\mathrm{d}\phi$). Given a Hamiltonian function $f$ on $\mathbb{S}^{2}$, the corresponding symplectic vector field $X_{f}$ is defined by
@@
Thus, the fluid velocity field is the symplectic vector field for the Hamiltonian given by the stream function: $\mathbf{v}=X_{\psi}$.
The space $\mathfrak{X}_{\mu}(\mathbb{S}^{2})$ of smooth, symplectic vector fields on $\mathbb{S}^{2}$ forms an infinite-dimensional Lie algebra (with the vector field bracket $[\cdot,\cdot]_{\mathfrak{X}}$). This algebra is isomorphic to the Poisson algebra of smooth functions modulo constants $C^{\infty}(\mathbb{S}^{2})/\mathbb{R}$ via the mapping $\psi\mapsto X_{\psi}$. Indeed, we have that
$-[X_{\psi},X_{\xi}]_{\mathfrak{X}}=\{\psi,\xi\}=\nabla^{\bot}\psi\cdot\nabla\xi.$
@@
Let $\mathfrak{g}$ be a Lie algebra. Its dual $\mathfrak{g}^{*}$ is a Poisson manifold via its Lie–Poisson bracket. In the infinite-dimensional case of smooth Hamiltonian (or stream) functions, i.e., $\mathfrak{g}=C^{\infty}(\mathbb{S}^{2})/\mathbb{R}$, one usually works with the smooth dual, which is constructed so that $C^{\infty}(\mathbb{S}^{2})^{*}\simeq C^{\infty}(\mathbb{S}^{2})$ via the $L^{2}$-pairing $\langle\omega,\psi\rangle=\int_{\mathbb{S}^{2}}\omega\psi\mu$. As indicated, we think of vorticity $\omega$ as a dual variable to the stream function $\psi\in\mathfrak{g}=C^{\infty}(\mathbb{S}^{2})/\mathbb{R}$. Thus,
@@
where $\dot{\omega}$ denotes time differentiation and $\operatorname{ad}^{*}_{f}\colon\mathfrak{g}^{*}\to\mathfrak{g}^{*}$ is defined by $\langle\operatorname{ad}^{*}_{f}\omega,\xi\rangle=\langle\omega,\{f,\xi\}\rangle$ for all $\xi\in\mathfrak{g}$. In the case $\mathfrak{g}=C^{\infty}(\mathbb{S}^{2})/\mathbb{R}$, with smooth dual $\mathfrak{g}^{*}=C^{\infty}_{0}(\mathbb{S}^{2})$, we obtain from the divergence theorem that
$\langle\omega,\{f,\xi\}\rangle=\int_{\mathbb{S}^{2}}\omega\nabla^{\bot}f\cdot\nabla\xi\,\mu=\langle-\operatorname{div}(\omega\nabla^{\bot}f),\xi\rangle=\langle\underbrace{-\{f,\omega\}}_{\operatorname{ad}^{*}_{f}\omega},\xi\rangle.$
Hence, $\operatorname{ad}^{*}_{f}\omega$ is minus the Poisson bracket, which reflects that the $L^{2}$-pairing on the Poisson space $C^{\infty}(\mathbb{S}^{2})$ is bi-invariant.
@@
$H(\omega)=\frac{1}{2}\int_{\mathbb{S}^{2}}\left|\nabla\psi\right|^{2}\,\mu=-\frac{1}{2}\int_{\mathbb{S}^{2}}\omega\underbrace{\Delta^{-1}\omega}_{\psi}\,\mu.$
Its variational derivative is $\frac{\delta H}{\delta\omega}=-\psi$. Consequently, the Lie–Poisson form (4) of equation (2) is
$\dot{\omega}+\{\psi,\omega\}=0,\quad\Delta\psi=\omega,$ (5)
@@
From (5) it is clear that the vorticity function $\omega$ is infinitesimally transported by the time-dependent Hamiltonian vector field $\mathbf{v}=X_{\psi}$. Consequently, the flow map $\Phi\colon\mathbb{S}^{2}\times\mathbb{R}\to\mathbb{S}^{2}$ obtained by integrating the vector field
$\frac{\partial{\Phi(x,t)}}{\partial t}=\mathbf{v}\big{(}\Phi(x,t),t\big{)},\quad\Phi(x,0)=x,$
is a curve $t\mapsto\Phi(\cdot,t)$ in the Lie group $G$ corresponding to the Lie algebra $\mathfrak{g}$. Since the integration of a vector field yields a symplectic map if and only if the vector field is symplectic, it follows that the Lie group consists of symplectic diffeomorphisms of $\mathbb{S}^{2}$, denoted $\operatorname{Diff}_{\mu}(\mathbb{S}^{2})$. Since $\omega$ is transported, we have that $\omega(x,t)=\omega(\Phi^{-1}(x,t),0)$, or in more compact notation $\omega_{t}=\omega_{0}\circ\Phi_{t}^{-1}$. The direct implication is that $\omega$ remains on the same co-adjoint orbit as $\omega_{0}$ (the orbits for the action of $\operatorname{Diff}_{\mu}(\mathbb{S}^{2})$ on $C^{\infty}_{0}(\mathbb{S}^{2})$). Therefore, since the map $\Phi_{t}$ has unitary Jacobian determinant, any functional of the form
$C_{f}(\omega)=\frac{1}{2\pi}\int_{\mathbb{S}^{2}}f\big{(}\omega(x,t)\big{)}\,\mu(x),\quad f\colon\mathbb{R}\to\mathbb{R},$

### v3

@@ Abstract.
Two-dimensional (2-D) incompressible, inviscid fluids produce fascinating patterns of swirling motion. How and why the patterns emerge are long-standing questions, first addressed in the 19th century by Helmholtz, Kirchhoff, and Kelvin. Countless researchers have since contributed to innovative techniques and results. Yet, the overarching problem of swirling 2-D motion and its long-time behavior remains largely open. Here we shed light on this problem via a link to isospectral matrix flows. The link is established through V. Zeitlin’s beautiful model for the numerical discretization of Euler’s equations in 2-D. When considered on the sphere, Zeitlin’s model offers deep connections between 2-D hydrodynamics and unitary representations of the rotation group. Consequently, it provides a dictionary that maps hydrodynamical concepts to matrix Lie theory, which in turn gives connections to matrix factorizations, random matrices, and integrability theory, for example. Results about finite-dimensional matrices can then be transferred to infinite-dimensional fluids via quantization theory, which is here used as an analysis tool (albeit traditionally describing the limit between quantum and classical physics). We demonstrate how the dictionary is constructed and how it unveils techniques for 2-D hydrodynamics. We also give accompanying convergence results for Zeitlin’s model on the sphere.
@@ 1. Introduction
If you stir up an incompressible, low-viscosity fluid confined to a thin (almost 2-D) bounded domain then, under conditions on the initial configuration and the applied forcing, what you see is spectacular, completely different from the situation in full 3-D. The fluid self-organizes into large, coherently swirling vortex condensates. [^1] The effect can be experimentally reproduced by a thin soap film flowing rapidly through a fine comb [13], or by a conducting fluid confined to a thin layer and driven into turbulence by a temporally varying magnetic field [51]. Some regions swirl counterclockwise (positive vorticity) while others swirl clockwise (negative vorticity), with occasional merger between equal-signed regions, and repulsion between opposite-signed regions. What is the mechanism of vortex condensation and under which conditions does it occur? More mathematically formulated, what is the generic long-time behavior of the 2-D Euler equations?
@@
A milestone was reached in 1949, when Onsager [48] suggested to use statistical mechanics applied to a large but finite number $n$ of point vortices on a flat torus (doubly periodic square). Onsager ingeniously realized that if the energy of a configuration is large enough, the thermodynamical temperature is negative. Point vortices of equal sign will then tend to cluster “so as to use up excess energy at the least possible cost in terms of degrees of freedom” [48, p. 281]. Thus, statistical mechanics predicts vortex condensation. Under the ergodicity assumption, it also predicts the long-time states corresponding to “thermal equilibrium”. Onsager’s ideas have fostered several mathematically rigorous results, for example weak convergence, in both positive and negative temperature regimes, of the canonical Gibbs measure as $n\to\infty$ [7, 34].
But there is a problem with the point vortex approach, as Onsager himself pointed out: “When we compare our idealized model with reality, we have to admit one profound difference: the distribution of vorticity which occur in the actual flow of normal liquids are continuous” [48, p. 281]. In particular, solutions to the 2-D Euler equations with continuous, bounded vorticity possess a rich geometric structure: they are described by an infinite-dimensional Lie–Poisson system, which gives rise to infinitely many conservation laws given by Casimir functions (cf. Sec. 2 below). Whereas point-vortices in principle adhere to this structure, with Casimirs corresponding to point vortex strengths, they do it in a weak sense which fails to capture the effects of the incompressibility for bounded vortices, for example. Furthermore, there is evidence that the continuous Casimirs significantly affect the long-time behavior (cf. Abramov and Majda [1]). A further criticism of Onsager’s statistical model is that the ergodicity assumption is not valid (which does not, however, wholly invalidate the statistical theory; see the discussion in Marchioro and Pulvirenti [42, Sec. 7.5]). From these perspectives, Onsager’s model is inadequate.
@@
Another established technique for 2-D hydrodynamics is non-linear PDE analysis. It has provided many rigorous and deep results, but at the cost of restricted settings, for example perturbations of steady solutions. We refer to the monograph of Marchioro and Pulvirenti [42] and the lecture notes of Šverák [53] for an overview. The grand vision, of course, is to bridge PDE analysis results with predictions from statistical hydrodynamics. It is a notoriously difficult problem. A flavor of the difficulty is captured in the following “incompatibility” between the two approaches: The statistically predicted vortex condensation implies merging of equally signed vorticity regions. Yet, in the 2-D Euler equations, vorticity is advected by the fluid velocity field, so merging can only occur by trapping increasingly narrow, but deep, vorticity variations in an overwhelmingly complicated stretch-and-fold process. These variations cannot “disappear” in the $C^{1}$-topology. Indeed, the $C^{1}$-norm along regular solutions can, and in numerical simulations typically does, grow extremely fast (e.g., double exponential [37]).
@@
A third tool for 2-D hydrodynamics is to carry out numerical experiments (e.g., Fig. 1, where the mixing phase (b) captures a typical stretch-and-fold process). Indeed, since the work of Lorentz [41] in the 1960s, computer simulations have guided many analytical results in fluid dynamics; recent examples are found in the references [14, 50, 37, 20, 9, 57]. But how can we know if the numerical results are consistent with the geometric structures principal to the long-time behavior? One approach is to ensure that the finite-dimensional dynamical model resulting from the discretization is itself a Lie–Poisson system, and that the finite-dimensional Lie–Poisson bracket approximates the infinite-dimensional one. Such a model was provided by Zeitlin [60, 61], based on quantization results of Hoppe [27, 28]. It is the only known Lie–Poisson spatial discretization of 2-D Euler, and numerical simulations indicate that it captures the qualitative behavior more accurately than traditional discretizations (e.g., the predicted spectral scaling law, which standard methods fail to catch even at significantly higher resolutions [10]). Yet, the model is surprisingly underexplored.
@@
Our paper draws from the following observation: in addition to numerical advantages, matrix hydrodynamics – the isospectral matrix flows enabled via Zeitlin’s model – brings theoretical insights to classical 2-D hydrodynamics, as it provides a link to matrix theory with all its findings. We therefore advocate matrix hydrodynamics as a tool for 2-D turbulence, similar in spirit to the point vortex model of Onsager, but with the benefit that it captures the Lie–Poisson geometry of continuous vorticity fields, including conservation of continuous Casimirs.
In addition to the examples in this paper, a motivation and feasibility check for the approach comes from complex geometry. Indeed, in the late 1980s, quantization was promoted to address problems related to Kähler–Einstein metrics and the Calabi conjecture. Tian [54] then gave convergence results, since used regularly in theoretical contexts. Only much later did Donaldson [17] use quantization in a numerical study of Kähler–Einstein metrics. This is the antithesis to the developments in 2-D hydrodynamics, where Zeitlin’s model has been a numerical tool, until recently.
@@
The primary aim of the paper is to give basic convergence results for matrix hydrodynamics in the spherical setting (Theorem 4, Theorem 6, Corollary 17, and Theorem 19), and to showcase the notion in selected examples and relate it to existing theories (Sec. 4-6). The secondary aim, addressed in Sec. 3, is to give a transparent derivation of Zeitlin’s model on the sphere, based on unitary representation theory and focused on geometry rather than algebraic formulae. Before that, we review in Sec. 2 the geometric structure of the 2-D Euler equations.
Acknowledgements This work was supported by the Swedish Research Council (grant number 2022-03453), the Knut and Alice Wallenberg Foundation (grant numbers WAF2019.0201 and KAW2020.0287), and the Göran Gustafsson Foundation for Research in Natural Sciences and Medicine. The computations were enabled by resources provided by Chalmers e-Commons at Chalmers. The data handling was enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725. The authors would like to thank Michele Dolce and Theo Drivas for fruitful discussions related to this work, and Michael Roop for important comments related to the convergence proof in Appendix A.
@@ 2. Background: geometry of the 2-D Euler equations
For an incompressible, inviscid fluid, Euler [21] showed how Newton’s equations give rise to a set of partial differential equations (PDEs) for the velocity field $\boldsymbol{v}(x,t)$ of the motion,
$\frac{\partial\boldsymbol{v}}{\partial t}+\nabla_{\boldsymbol{v}}\boldsymbol{v}=-\nabla p,\quad\operatorname{div}\boldsymbol{v}=0,$ (1)
where $p$ is the pressure function and $\nabla_{\boldsymbol{v}}$ is the co-variant derivative (furthermore, $\boldsymbol{v}$ is tangential to the boundary if present). Euler’s equations (1) make sense on any Riemannian manifold, but in 2-D the geometric structure is richer than in higher dimensions, and the dynamics is significantly different. Throughout this paper we shall work on the sphere $\mathbb{S}^{2}$, partly because of its relevance in geophysical contexts, but mostly because it later allows a description in terms of unitary representation theory that renders the geometric structures in Zeitlin’s model more transparent.
Let $J\colon T\mathbb{S}^{2}\to T\mathbb{S}^{2}$ denote the bundle mapping for rotation by $\pi/2$ in positive direction. The curl operator maps the vector field $\boldsymbol{v}$ on $\mathbb{S}^{2}$ to the vorticity function
$\omega=\operatorname{curl}\boldsymbol{v}\equiv\operatorname{div}(J\boldsymbol{v}).$
@@
$\frac{\partial\omega}{\partial t}+\operatorname{div}(\omega\boldsymbol{v})=0.$
Further, since the first co-homology of the sphere is trivial, it follows from Hodge theory that the divergence free vector field $\boldsymbol{v}$ can be written $\boldsymbol{v}=\nabla^{\bot}\psi\equiv-J\nabla\psi$, where $\psi$ is the stream function, unique up to a constant. Euler’s equations (1) is then formulated entirely in terms of the vorticity and stream functions
$\frac{\partial\omega}{\partial t}+\operatorname{div}(\omega\nabla^{\bot}\psi)=0,\quad-\Delta\psi=\omega.$ (2)
Following Arnold [2] and Marsden and Weinstein [43], we now demonstrate how to arrive at these equations from the point-of-view of symplectic geometry and Hamiltonian dynamics.
The sphere is a symplectic manifold for the symplectic form given by the spherical area form $\mu$ (in spherical coordinates $(\theta,\phi)\in[0,\pi]\times[0,2\pi)$ the area form is $\mu=\sin\theta\,\mathrm{d}\theta\wedge\mathrm{d}\phi$). Given a Hamiltonian function $f$ on $\mathbb{S}^{2}$, the corresponding symplectic vector field $X_{f}$ is defined by
@@
Thus, the fluid velocity field is the symplectic vector field for the Hamiltonian given by the stream function: $\boldsymbol{v}=X_{\psi}$.
The space $\mathfrak{X}_{\mu}(\mathbb{S}^{2})$ of smooth, symplectic vector fields on $\mathbb{S}^{2}$ forms an infinite-dimensional Lie algebra (with the vector field bracket $\cdot,\cdot]_{\mathfrak{X}}$). This algebra is isomorphic to the Poisson algebra of smooth functions modulo constants $C^{\infty}(\mathbb{S}^{2})/\mathbb{R}$ via the mapping $\psi\mapsto X_{\psi}$. Indeed, we have that
$X_{\psi},X_{\xi}]_{\mathfrak{X}}=X_{\{\psi,\xi\}}\qquad\text{where}\qquad\{\psi,\xi\}=\nabla\psi\cdot\nabla^{\bot}\xi.$
@@
Let $\mathfrak{g}$ be a Lie algebra. Its dual $\mathfrak{g}^{*}$ is a Poisson manifold via its Lie–Poisson bracket defined on functions $F,H\in C^{\infty}(\mathfrak{g}^{*})$ by
$\prec\!F,H\!\succ\!(\omega)=\langle\omega,[\mathrm{d}F(\omega),\mathrm{d}H(\omega)]\rangle,\qquad\omega\in\mathfrak{g}^{*},$ (4)
where $\cdot$$\cdot$] denotes the Lie bracket on $\mathfrak{g}$ and $\langle\cdot,\cdot\rangle$ is the pairing between $\mathfrak{g}^{*}$ and $\mathfrak{g}$. The Lie-Poisson bracket (4) originates from the canonical symplectic structure on the co-tangent bundle of the corresponding Lie group $G$, via symmetry reduction $T^{*}G/G\simeq\mathfrak{g}^{*}$. For details on this framework and its application in hydrodynamics, we refer to Arnold and Khesin [3].
In the infinite-dimensional case of the group of symplectomorphisms on $\mathbb{S}^{2}$ and its Lie algebra of smooth Hamiltonian (or stream) functions, i.e., $G=\mathrm{Diff}_{\mu}(\mathbb{S}^{2})$ and $\mathfrak{g}=C^{\infty}(\mathbb{S}^{2})/\mathbb{R}$, one usually restricts to the smooth dual, which is constructed so that $C^{\infty}(\mathbb{S}^{2})^{*}\simeq C^{\infty}(\mathbb{S}^{2})$ via the $L^{2}$-pairing $\langle\omega,\psi\rangle=\int_{\mathbb{S}^{2}}\omega\psi\mu$. As indicated, we think of vorticity $\omega$ as a dual variable to the stream function $\psi\in\mathfrak{g}=C^{\infty}(\mathbb{S}^{2})/\mathbb{R}$. Thus,
@@
where $\dot{\omega}$ denotes time differentiation and $\operatorname{ad}^{*}_{\psi}\colon\mathfrak{g}^{*}\to\mathfrak{g}^{*}$ is defined by $\langle\operatorname{ad}^{*}_{\psi}\omega,\xi\rangle=\langle\omega,\{\psi,\xi\}\rangle$ for all $\xi\in\mathfrak{g}$. In the case $\mathfrak{g}=C^{\infty}(\mathbb{S}^{2})/\mathbb{R}$, with smooth dual $\mathfrak{g}^{*}=C^{\infty}_{0}(\mathbb{S}^{2})$, we obtain from the divergence theorem that
$\langle\omega,\{\psi,\xi\}\rangle=-\int_{\mathbb{S}^{2}}\omega\nabla^{\bot}\psi\cdot\nabla\xi\,\mu=\langle\operatorname{div}(\omega\nabla^{\bot}\psi),\xi\rangle=\langle\underbrace{\{\omega,\psi\}}_{\operatorname{ad}^{*}_{\psi}\omega},\xi\rangle.$
Hence, $\operatorname{ad}^{*}_{\psi}\omega$ is minus the Poisson bracket, which reflects that the $L^{2}$-pairing on the Poisson space $C^{\infty}(\mathbb{S}^{2})$ is bi-invariant.
@@
$H(\omega)=\frac{1}{2}\int_{\mathbb{S}^{2}}\left|\nabla\psi\right|^{2}\,\mu=\frac{1}{2}\int_{\mathbb{S}^{2}}\omega\underbrace{(-\Delta)^{-1}\omega}_{\psi}\,\mu.$
Its variational derivative is $\frac{\delta H}{\delta\omega}=\psi$. Consequently, the Lie–Poisson form (5) of equation (2) is
$\dot{\omega}+\{\omega,\psi\}=0,\quad-\Delta\psi=\omega,$ (6)
@@
From (6) it is clear that the vorticity function $\omega$ is infinitesimally transported by the time-dependent Hamiltonian vector field $\boldsymbol{v}=X_{\psi}$. Consequently, the flow map $\Phi\colon\mathbb{S}^{2}\times\mathbb{R}\to\mathbb{S}^{2}$ obtained by integrating the vector field
$\frac{\partial{\Phi(x,t)}}{\partial t}=\boldsymbol{v}\big{(}\Phi(x,t),t\big{)},\quad\Phi(x,0)=x,$
is a curve $t\mapsto\Phi(\cdot,t)$ in the Lie group $G$ corresponding to the Lie algebra $\mathfrak{g}$. Since the integration of a vector field yields a symplectic map if and only if the vector field is symplectic, it follows that the Lie group consists of symplectic diffeomorphisms of $\mathbb{S}^{2}$, denoted $\operatorname{Diff}_{\mu}(\mathbb{S}^{2})$. Since $\omega$ is transported, we have that $\omega(x,t)=\omega(\Phi^{-1}(x,t),0)$, or in more compact notation $\omega_{t}=\omega_{0}\circ\Phi_{t}^{-1}$. The direct implication is that $\omega$ remains on the same coadjoint orbit as $\omega_{0}$ (the orbits for the action of $\operatorname{Diff}_{\mu}(\mathbb{S}^{2})$ on $C^{\infty}_{0}(\mathbb{S}^{2})$). Therefore, since the map $\Phi_{t}$ has unitary Jacobian determinant, any functional of the form
$C_{f}(\omega)=\int_{\mathbb{S}^{2}}f\big{(}\omega(x,t)\big{)}\,\mu(x),\quad f\colon\mathbb{R}\to\mathbb{R}$,